Malko Enterprises’ bonds currently sell for $1,150. They have a 6-year maturity, an annual coupon of $75, and a par value of $1,000. What is their current yield?

a. 5.74% b. 7.30% c. 5.22% d. 7.76% e. 6.52%
Answer #1

formula for current yield = coupon rate / price

here, price = 1150

coupon rate = 75

therefore, 75 / 1150 = 0.065217

= 0.065217 * 100 = 6.52%

so current yield is option e, 6.52%